OpenStudy (anonymous):

Please Simplify this for me (x^½)³×(√x)^9

OpenStudy (pradipgr817):

x^(3/2) x x^(9/2) = x^(3/2 + 9/2)=x^6

OpenStudy (anonymous):

x^3*(1/8)*9x^(1/2)